OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale?

OpenStudy (anonymous):

Just substitute 100 in place of x in the given equation

OpenStudy (bahrom7893):

what Harkirat said: C = 0.765*100 + 0.06(0.765*100) = 76.5 + 0.06*76.5

OpenStudy (bahrom7893):

u can just finish the rest on a calculator